Two charges +3 μC and +12 μC are fixed 1 m apart, with the second one to the right. Find the magnitude and direction of the net force on a −2nC charge when placed at the following locations: (a) halfway between the two (b) half a meter to the left of the +3μC charge (c) half a meter above the +12μC charge in a direction perpendicular to the line joining the two fixed charges.


Answers:

a. Blank 1 uN

b. Blank 2 uN

c. Blank 3i + Blank 4j uN


1
Expert's answer
2021-05-09T13:07:39-0400

(a) Let "q_1=3\ \mu C", "q_2=-2\ nC" and "q_3=12\ \mu C".Then, we can find the net force on a −2nC charge as follows:


"F_{net}=F_{32}-F_{12},""F_{net}=\dfrac{k}{r^2}(|q_2q_3|-|q_1q_2|),"

"F_{net}=\dfrac{8.99\cdot10^9\ \dfrac{Nm^2}{C^2}}{(0.5\ m)^2}(|(-2\cdot10^{-9}\ C)\cdot12\cdot10^{-6}\ C|-|3\cdot10^{-6}\ C\cdot(-2\cdot10^{-9}\ C)|)=6.47\cdot10^{-4}\ N."

The sign plus means that the net force directed to the right.

(b) Let "q_1=-2\ nC", "q_2=3\ \mu C" and "q_3=12\ \mu C".Then, we can find the net force on a −2nC charge as follows:


"F_{net}=F_{21}+F_{31},""F_{net}=k(\dfrac{|q_1q_2|}{r_{21}^2}+\dfrac{|q_1q_3|}{r_{31}^2}),"

"F_{net}=8.99\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(\dfrac{|(-2\cdot10^{-9}\ C)\cdot3\cdot10^{-6}\ C|}{(0.5\ m)^2}+\dfrac{|(-2\cdot10^{-9}\ C)\cdot12\cdot10^{-6}\ C|}{(1.5\ m)^2})=3.12\cdot10^{-4}\ N."

The sign plus means that the net force directed to the right.

(c) Let "q_1=3\ \mu C", "q_2=12\ \mu C" and "q_3=-2\ nC". Let's first find the "x"- and "y"-components of the electric force:


"F_{13,x}=-k\cdot\dfrac{|q_1q_3|}{r_{13}^2}\cdot\dfrac{r_{23}}{r_{13}},"

"F_{13,x}=-8.99\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot\dfrac{|3\cdot10^{-6}\ C\cdot(-2\cdot10^{-9}\ C)|}{(\sqrt{(1\ m)^2+(0.5\ m)^2})^2}\cdot\dfrac{0.5\ m}{\sqrt{(1\ m)^2+(0.5\ m)^2}}=-1.93\cdot10^{-5}\ N,"


"F_{23,x}=0\ N,""F_{13,y}=-k\cdot\dfrac{|q_1q_3|}{r_{13}^2}\cdot\dfrac{r_{12}}{r_{13}},"

"F_{13,y}=-8.99\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot\dfrac{|3\cdot10^{-6}\ C\cdot(-2\cdot10^{-9}\ C)|}{(\sqrt{(1\ m)^2+(0.5\ m)^2})^2}\cdot\dfrac{1.0\ m}{\sqrt{(1\ m)^2+(0.5\ m)^2}}=-3.86\cdot10^{-5}\ N,"

"F_{23,y}=-k\dfrac{|q_2q_3|}{r_{23}^2},"

"F_{23,y}=-8.99\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot\dfrac{|12\cdot10^{-6}\ C\cdot(-2\cdot10^{-9}\ C)|}{(0.5\ m)^2}=-8.63\cdot10^{-4}\ N,"


"F_x=F_{13,x}=-1.93\cdot10^{-5}\ N,""F_y=F_{13,y}+F_{23,y},""F_y=-3.86\cdot10^{-5}\ N+(-8.63\cdot10^{-4}\ N)=-9.02\cdot10^{-4}\ N."

We can find the net electric force on "q_3" from the Pythagorean theorem:


"F=\sqrt{F_x^2+F_y^2},""F=\sqrt{(-1.93\cdot10^{-5}\ N)^2+(-9.02\cdot10^{-4}\ N)^2}=9.02\cdot10^{-4}\ N."

We can find the direction of the net electric force from the geometry:


"\theta=cos^{-1}\dfrac{F_x}{F},""\theta=cos^{-1}\dfrac{-1.93\cdot10^{-5}\ N}{9.02\cdot10^{-4}\ N}=91.2^{\circ}."

The net electric force on charge "q_3"directed "91.2^{\circ}" below "x"-axis.